OpenStudy (anonymous):

What is the solution to the equation 36^3c=sqrt6^(8c+6)

OpenStudy (anonymous):

I think it's 3/2?

OpenStudy (anonymous):

\[36^{3c}=(\sqrt{6})^{8c+6} \\ then \ \ \ (6^2)^{3c}=(6^{1/2})^{8c+6} \\ and \ \ \ (6)^{6c}=(6)^{4c+3} \\ so \ \ \ 6c=4c+3 \\ 2c=3 \\ c=3/2 \]

OpenStudy (anonymous):

line 2 : (a^b)^c=a^bc